OpenStudy (anonymous):

Find the indefinite integral of (x^2/(sqrt(1-x))).

OpenStudy (anonymous):

\[\int\limits \frac{ x^2 }{ \sqrt{1-x} } dx\]

OpenStudy (helder_edwin):

u can try integration by parts